OpenStudy (anonymous):

Find the exact value of tan (arcsin(2/5)). Please explain how you got your answer

OpenStudy (sirm3d):

\[\arcsin (2/5)\] means the angle (theta) such that \[\sin \theta = 2/5\] drawing the right triangle 

OpenStudy (sirm3d):

find the length of the third side, from which you can compute the value of tan(theta)

OpenStudy (anonymous):

i don reamer how to do that, is there a equation for that

OpenStudy (anonymous):

im sorry i dont remember

OpenStudy (anonymous):

is there a equation for that???????

OpenStudy (sirm3d):

LET \[\large \theta =\arcsin \frac{ 2 }{ 5 }\] the equation is the same as \[\large \sin \theta=\frac{ 2 }{ 5 }\]

OpenStudy (sirm3d):

the expression \[\large \tan(\arcsin \frac{ 2 }{ 5 })\] is now changed to \[\large \tan \theta\] a value you must find from the condition \[\large \sin \theta = \frac{ 2 }{ 5 }\]

OpenStudy (anonymous):

so is it 0.411 +/- 2 pi n

OpenStudy (anonymous):

or 2.7301 +/- 2 pi n